OpenStudy (anonymous):

Medal You bought a car for $6000 down and made payments of $299.50 each month for 36 months. (a) Find the amount of the payments over the 36 months. $ ? (b) Find the total cost of the car. $ ?

OpenStudy (texaschic101):

amount of payments : 299.50 * 36 = your answer total cost : 6000 + (299.50 * 36) = total cost you can do the math

OpenStudy (anonymous):

16782

OpenStudy (anonymous):

am I right?

OpenStudy (anonymous):

For a) What is 299.50 * 36 ?

OpenStudy (anonymous):

Do it on your calculator, if you're allowed to.

OpenStudy (anonymous):

10,782

OpenStudy (anonymous):

So that's your answer for a --- 10, 782

OpenStudy (anonymous):

To get the answer for b ---- 6,000 + 10,782 = *your answer*

OpenStudy (anonymous):

16,782

OpenStudy (anonymous):

Exactly a~ 10, 782 b~ 16,782

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

You're Welcome :) Just tag me anytime if you need help, and I'll see what I can do.

OpenStudy (anonymous):

ok :)